# Graph neural network surrogate for strategic transport planning

### v1

@@ Abstract
As the complexities of urban environments continue to grow, the modelling of transportation systems become increasingly challenging. This paper explores the application of advanced Graph Neural Network (GNN) architectures as surrogate models for strategic transport planning. Building upon a prior work that laid the foundation with graph convolution networks (GCN), our study delves into the comparative analysis of established GCN with the more expressive Graph Attention Network (GAT). Additionally, we propose a novel GAT variant (namely GATv3) to address over-smoothing issues in graph-based models. Our investigation also includes the exploration of a hybrid model combining both GCN and GAT architectures, aiming to investigate the performance of the mixture. The three models are applied to various experiments to understand their limits. We analyse hierarchical regression setups, combining classification and regression tasks, and introduce fine-grained classification with a proposal of a method to convert outputs to precise values. Results reveal the superior performance of the new GAT in classification tasks. To the best of the authors’ knowledge, this is the first GAT model in literature to achieve larger depths. Surprisingly, the fine-grained classification task demonstrates the GCN’s unexpected dominance with additional training data. This shows that synthetic data generators can increase the training data, without overfitting issues whilst improving model performance. In conclusion, this research advances GNN based surrogate modelling, providing insights for refining GNN architectures. The findings open avenues for investigating the potential of the newly proposed GAT architecture and the modelling setups for other transportation problems.
@@
journal: a journal
@@ 1 Introduction
Transportation plays a critical role in today’s society. The modern urban landscape is characterized by ever-increasing mobility demands, necessitating sophisticated and efficient transport systems. Strategic transport planning plays a pivotal role in shaping the future, to meet the needs of growing populations and evolving societal demands. However, the complexity of urban environments, with their intricate networks of roads, public transit systems, and diverse travel patterns, presents a formidable challenge in accurately modelling and simulating transportation systems.
@@
In recent years, the development of surrogate models has emerged as a promising approach to tackle the complexity of strategic transport planning. Surrogate models are easier to apply and have been shown to even surpass the performance of the original models in other fields (e.g., Choi & Kumar, 2024). Furthermore, such models can also be tuned directly using data and auto-calibration procedures are available. User-friendly models like these that can evaluate complex mobility options are being expected much by the city authorities [Narayanan et al., 2023].
For strategic transport planning, a prior study [Narayanan et al., 2024] laid the foundation by investigating the feasibility of constructing a surrogate through the application of Graph Convolution Network (GCN) in both classification and regression setups. This is a primary work to ascertain whether Graph Neural Network (GNN) based surrogates can even replicate the traditional strategic modelling approach (i.e., the four step model), which could pave way to replicate complex approaches like agent-based models in future. While the study demonstrated initial promise, the potential for further refinement and enhancement of surrogate modelling techniques remains untapped.
This paper builds upon the foundation laid in the pioneer work of Narayanan et al. [2024], through the exploration of advanced neural network architectures and different task setups. We delve into the comparative analysis of the established GCN model with the more recent and expressive Graph Attention Network (GAT). Additionally, we introduce a novel variant of GAT tailored to address the prevalent issue of over-smoothing in graph-based models. Furthermore, to push the boundaries of surrogate modelling even further, we investigate the synergistic potential of combining both GCN and GAT architectures. We also explore the realm of hierarchical regression setups, where we explore the integration of classification and regression tasks. This holistic approach seeks to investigate whether there exist benefits in combining both the types of tasks.
To avoid the issues associated with the complexity of regression and hierarchical setups while ensuring the possibility of obtaining precise values, we delve into the realm of fine-grained classification combined with a method to convert its output to precise values. Finally, a synthetic data generator is implemented to study the impact of increasing the training data. Through this research endeavor, we aim to not only expand the horizons of GNN based surrogate modelling in strategic transport planning, but also provide insights and methodologies for enhancing the applicability and usability of such models in urban planning scenarios.
The methodological framework of this study is outlined in Figure 1. Section 2 provides a brief review of the prior work. Section 3 elaborates on the methodology, detailing the datasets and the architectures of GAT and the proposed variant. It also introduces the hybrid model (GCN+GAT). Section 5 presents the experimental results, which consists of (i) comparison between the different GNN models, (ii) hierarchical regression task and (iii) fine grained classification task. Finally, Section 6 offers concluding remarks and outlines avenues for future research.
Figure 1: Methodological framework
@@ 2.1 GNNs for transport planning
GNNs are slowly gaining significant traction in the domain of transport planning due to their ability to capture intricate spatial dependencies within complex urban networks. For example, GNNs are used to predict the demand for micromobility modes. In this direction, Lin et al. [2018] proposes Graph Convolutional Neural (GCN) Network model to forecast station-level hourly demand for a bike-sharing system in New York City and conclude that such models show promising results. Ding et al. [2022] utilised a GNN model to study the effect of a dockless bike-sharing scheme on the the demand of the London Cycle Hire scheme. For a shared E-scooter system, Song et al. [2023] developed a GNN model to predict the demand in Louisville, USA. Both of these studies also conclude that GNN models make better predictions.
GNNs are also used for studying taxi trips. For example, Wang et al. [2024] developed a spatio-temporal GNN model for taxi route assignment. Tygesen et al. [2023] conduct experiments on NYC Yellow Taxi and provide insights on the kinds of connections GNNs use for spatio-temporal predictions in the transport domain. They also perform tests for short term traffic prediction. Similarly, Cui et al. [2020] incorporate graph wavelet and employ gated recurrent structure for network-wide traffic forecasting. Xu et al. [2023] conclude that graph based models perform better than other machine learning models for network-wide short-term traffic speed prediction and imputation. Graph based neural network models are also used for real-time incident prediction (e.g., Tran et al., 2023) and can be helpful in solving the air transportation nexus among safety, efficiency and resilience [Wandelt et al., 2024].
@@ 2.2 Surrogate modelling in the field of transportation
Surrogate modelling is a macro-modeling technique to approximate expensive functions and to replicate input–output relationships. Surrogate models are also called meta- or response surface models and posses beneficial qualities, such as cheaper evaluation costs and sometimes, better accuracy. To approximate complex simulation systems, deep learning based surrogate models are appropriate, as they are suitable to represent non-linear relationships and can handle high dimensional problems well. In the field of traffic forecasting, Vlahogianni [2015] introduces the concept of surrogate modelling and suggests some approaches to specific optimization problems that a transportation researcher may face when dealing with forecasting models. The approaches can be adapted to various transportation problems and the results convey that surrogate models can generate more accurate predictions in significantly reduced times.
Focusing on dynamic user equilibrium, Ge et al. [2020] formulates a dynamic user equilibrium traffic assignment model incorporating surrogate dynamic network loading models. They conclude that the usage of surrogate approaches reduces the computational times by up to 90%, while at the same time, leading to lower errors. Surrogate models are also used to assign Dedicated Bus Lanes (DBLs) across a large-scale road network. The results of numerical studies from Li et al. [2022] show that surrogate based methods for optimal DBL allocation scheme can save a substantial amount of time, especially with limited computational resource, while also leading to around 5% better network performance than the costly simulation approaches.
The concept of GNN based surrogate modelling in the context of transport modelling emerged very recently, with the pioneering work of Narayanan et al. [2024]. They laid a foundational groundwork by investigating the feasibility of developing a GCN based surrogate for strategic transport planning. Their study delve into both classification and regression setups, and based their analysis by generating data from Greater Munich metropolitan region. The experimental results show that GNNs have the potential to act as transport planning surrogates and the deeper models perform better than their shallow counterparts. However, the models suffer performance degradation with an increase in network size.
@@ 2.3 Summary and research gaps
The study of Narayanan et al. [2024] demonstrated the efficacy of GCNs and showcased how GCNs outperformed traditional machine learning models by effectively leveraging the graph structure inherent in transportation networks. Nevertheless, the limitations of GCNs (e.g., the neighbourhood aggregation function used in the GCN puts equal weight on all neighbours, while all neighbours may not be equally relevant) and that of the classification and regression setups necessitate the investigation of the use of advanced GNN architectures and the analysis of novel modelling setups, which will be the focus of this study.
@@ 3 Methodological framework
This study builds over the methodological framework established in Narayanan et al. [2024]. Therefore, the procedural sequence is as follows: (i) data generation for the four-step model, (ii) simulation in the four-step model, (iii) data transformation (i.e., direct graph) and the creation of the training data (iv) GNN model training and model evaluation. The regularization methods (i.e., Dropout and GraphNorm), loss functions (cross entropy for the classification tasks and mean squared error for the regression tasks) and evaluation metrics (accuracy and F1-score for the classification tasks and mean absolute error and R2) from there are also utilised in this study. For detailed information of the setup, the reader is referred to the original study.
As an addition to the existing setup, this study develops and investigates another data generation process and further GNN models, which are introduced in the below sections. The implementation of the entire pipeline is performed in Python. The baseline models are estimated using Scikit-Learn [Pedregosa et al., 2011]. GNN models are implemented in PyTorch Geometric [Fey & Lenssen, 2019], which is an extension to the PyTorch library [Paszke et al., 2019]. Due to privacy concerns, the entire code (including the ones for data generation and processing) is provided in a private GitHub repository, which is found at https://github.com/nikita68/4_step_model_surrogate. Access to it is possible upon request. Nevertheless, the processed datasets are separately available at https://github.com/nikita6187/TransportPlanningDataset.
@@ 3.1 Synthetic data generation
For data generation, Narayanan et al. [2024] implemented augmented data generation, wherein the network of Munich metropolitan region is used to generate multiple subnetworks for model training. It is often seen that more data significantly helps deep learning models [Wang et al., 2017]. However, the original network data source utilised for the augmented dataset in Narayanan et al. [2024] has only a finite number of possible subnetworks, until the training dataset has a critical overlap with the validation and test datasets. To overcome this issue, while still providing arbitrarily more data, this study implements additional data generation procedure, namely the synthetic data generation.
A synthetic region generation algorithm is implemented, which does not use any original network as a source. This algorithm is motivated based on the works of [Parish & Müller, 2001, Smelik et al., 2014, Cogo et al., 2019]. Assuming that the target amount of nodes ($\mathds{G}_{nodes}$) is provided as input, along with the components for generating zones and zonal information (as identical to that of the augmented data generation procedure), the algorithm performs the following steps:
@@
The synthetic dataset is generated with 10000 samples, with a target number of nodes between 15 and 80. Note, the synthetic dataset is used in this paper only as supplementary training data for the augmented dataset with further details being provided in the experiments section (Section 4).
3.2 Graph attention networks
Narayanan et al. [2024] developed a GNN surrogate based on Graph Convolutions Networks (GCN). The neighbourhood aggregation function used in the GCN puts equal weight on all neighbours. However, not all neighbours may be equally relevant. An example would be an incoming road having a high traffic demand, whilst another having zero at a junction. Thus, the information relevant for the outgoing road comes more from the high demand incoming road. Therefore, in this section, the graph attention networks are presented, together with a novel extension to enable them to generalize to deeper networks.
Both GATv1 and GATv2 still suffer from the over-smoothing problem, much like the original GCN. This results in both models performing poorly when trying to stack more than 5 layers. To the best of the authors’ knowledge, there is no solution available to this problem at the time of this study. Inspired by the results of GCNII, a novel extension is proposed to use the same simple residual trick from [Chen et al., 2020]. Thus, the GATv2 formulation becomes as follows:
[…]
where,
$\alpha$ is a fixed hyperparameter
$\sigma$ is the activation function
$k$ represents all the direct neighbours of node $i$
$N_{ff}$ is a one layer MLP
[…]
[;][;][ ; ] operator concatenates the input vectors
Even though it is a minor change performed on the GATv2 model, this solves the issue of developing deeper GAT models, as will be shown in later sections. From now, this model is denoted as $GATv3$. Note that, due to the extra computations, the number of GAT layers which can fit onto one GPU is significantly less than those of GCNs.
3.3 Combined GAT and GCN
The GCNII implemented in Narayanan et al. [2024] and the GATv3 introduced in this study solve the over-smooting problem, but do not specifically focus on over-squashing. Over-squashing is the effect where each additional GNN layer increases exponentially the amount of information that needs to contained within the fixed sized feature vector of a node [Alon & Yahav, 2020]. The effect is shown to result in information loss, which is critical for proper modelling. To the best of the authors’ knowledge, there is no known adequate fix to this problem. The over-squashing problem can potentially pose a major problem for transport planning predictions, as information needs to be exchanged between nodes that are large distances away from each other. This is, especially, important for nodes containing zonal information.
To overcome the aformentioned issue, an attempt has been made to combine the GAT and GCN models. The idea is to first have a number of GATv3 layers, using an artificial graph connecting only zone nodes together. Next, the input is fed through a larger number of GCNII layers, which use the entire network graph. Finally, feed forward layers do the last processing of the data and output link level predictions. The intuition behind this model is similar to that of the strategic transport planning. The first GAT layers focus on determining the demand of the individual zones. The following GCN layers then focus on realistically applying the determined demand onto the existing network. The artificial graph used in the GAT layers is a graph where all zonal nodes are connected, whilst all other nodes are disconnected. This ensures a focus on the socio-economic information governing the demand aspect of transport planning. The model type is called the combined GAT and GCN model, and is abbreviated as $GCN+GAT$.
4 Experimental setup
This section will focus on three distinct setups, i.e., (i) a classification setup to compare the different GNNs, (ii) a hierarchical regression setup to investigate the effect of combining the classification and regression setups and (iii) a fine-grained classification setup to exploit the stability of classification, which also includes the implementation of a small transformation procedure to convert the classification results into actual traffic values. Finally, the impact of additional data will be explored using the fine-grained classification setup.
4.1 Comparison of GCNII, GATv2, GATv3 and GCN+GATv3

### v2

# Development of a graph neural network surrogate for travel demand modelling

@@ Abstract
As urban environments grow, the modelling of transportation systems becomes increasingly complex. This paper advances the field of travel demand modelling by introducing advanced Graph Neural Network (GNN) architectures as surrogate models, addressing key limitations of previous approaches. Building on prior work with Graph Convolutional Networks (GCNs), we introduce GATv3, a new Graph Attention Network (GAT) variant that mitigates over-smoothing through residual connections, enabling deeper and more expressive architectures. Additionally, we propose a fine-grained classification framework that improves predictive stability while achieving numerical precision comparable to regression, offering a more interpretable and efficient alternative. To enhance model performance, we develop a synthetic data generation strategy, which expands the augmented training dataset without overfitting. Our experiments demonstrate that GATv3 significantly improves classification performance, while the GCN model shows unexpected dominance in fine-grained classification when supplemented with additional training data. The results highlight the advantages of fine-grained classification over regression for travel demand modelling tasks and reveal new challenges in extending GAT-based architectures to complex transport scenarios. Notably, GATv3 appears well-suited for classification-based transportation applications, such as section control and congestion warning systems, which require a higher degree of differentiation among neighboring links. These findings contribute to refining GNN-based surrogates, offering new possibilities for applying GATv3 and fine-grained classification in broader transportation challenges.
@@
journal: Communications in Transportation Research
@@ 1 Introduction
Transportation plays a critical role in today’s society. The modern urban landscape is characterized by ever-increasing mobility demands, necessitating sophisticated and efficient transport systems. Focusing on long term planning, travel demand modelling plays a pivotal role in shaping the future, to meet the needs of growing populations and evolving societal demands. However, the complexity of urban environments, with their intricate networks of roads, public transit systems, and diverse travel patterns, presents a formidable challenge in accurately modelling and simulating transportation systems for the future. The need for scalable, adaptable, and computationally efficient modeling techniques has therefore become more pressing than ever.
@@
In recent years, the development of surrogate models has emerged as a promising approach to tackle the complexity of strategic transport planning. Unlike traditional models that often require extensive computational resources and expert calibration, surrogate models leverage data-driven methodologies to streamline the evaluation of complex mobility solutions. These models have demonstrated their potential in various domains by not only reducing computational costs but also enhancing predictive accuracy (e.g., Choi & Kumar, 2024). Moreover, such models can also be tuned directly using data and auto-calibration procedures are available. User-friendly models like these that can evaluate complex mobility options are being expected much by the city authorities [Narayanan et al., 2023].
A proof-of-concept by Narayanan et al. [2024] demonstrated the feasibility of using Graph Convolutional Networks (GCNs) as surrogates for travel demand modelling. This foundational work explored the potential of Graph Neural Networks (GNNs) in replicating the traditional four-step model, offering a pathway toward more sophisticated applications such as agent-based models. While these initial findings were promising, key challenges remain unaddressed - particularly concerning model expressiveness and stability in prediction accuracy. This paper builds upon and significantly advances the previous work by introducing novel methodologies that enhance both the depth and applicability of GNN-based surrogate modeling in strategic transport planning. Specifically, we introduce:
  • GATv3: A Graph Attention Network (GAT) variant that overcomes the depth limitations of existing GAT models caused by over-smoothing. By incorporating residual connections, GATv3 enables deeper and more expressive network architectures.
  • Synthetic data augmentation strategy: To address data scarcity and improve model performance, we introduce a synthetic data generator that effectively expands the augmented training dataset from Narayanan et al., 2024. Purely synthetic data is the one generated procedurally, whilst augmented implies that a true data source is employed, but modified to produce more data. This new approach is implemented to enrich the model’s learning capabilities without inducing overfitting.
  • Fine-grained classification framework: We propose an approach that transforms traditional regression problems into a fine-grained classification setup. This methodology not only ensures numerical precision comparable to regression but also enhances model stability, thereby making it more suitable for strategic transport applications.
The remainder of this paper is structured as follows: Section 2 provides a brief review of the prior work. Section 3 elaborates on the methodology, detailing the architectures of GAT and the proposed variant, synthetic data generation procedure and the different experimental setups. The experimental results are presented in Section 4. Finally, Section 5 offers concluding remarks.
@@ 2.1 GNNs for transport planning
GNNs are slowly gaining significant traction in the domain of transport planning due to their ability to capture intricate spatial dependencies within complex urban networks. For example, GNNs are used to predict the demand for micromobility modes. In this direction, Lin et al. [2018] proposed Graph Convolutional Neural (GCN) Network model to forecast station-level hourly demand for a bike-sharing system in New York City and conclude that such models show promising results. Ding et al. [2022] utilised a GNN model to study the effect of a dockless bike-sharing scheme on the the demand of the London Cycle Hire scheme. For a shared E-scooter system, Song et al. [2023] developed a GNN model to predict the demand in Louisville, USA. Both of these studies also conclude that GNN models make better predictions.
GNNs are also used for studying metro and taxi trips. For example, Zhao et al. [2023] designed an adaptive GCN model to predict short-term passenger flow for metro systems. Wang et al. [2024] developed a spatio-temporal GNN model for taxi route assignment. Tygesen et al. [2023] conduct experiments on NYC Yellow Taxi and provide insights on the kinds of connections GNNs use for spatio-temporal predictions in the transport domain. They also perform tests for short term traffic prediction. Similarly, Cui et al. [2020] incorporate graph wavelet and employ gated recurrent structure for network-wide traffic forecasting. Xu et al. [2023] conclude that graph based models perform better than other machine learning models for network-wide short-term traffic speed prediction and imputation. Graph based neural network models are also used for real-time incident prediction (e.g., Tran et al., 2023) and can be helpful in solving the air transportation nexus among safety, efficiency and resilience [Wandelt et al., 2024].
@@ 2.2 Surrogate modelling in the field of transportation
Surrogate modelling is a macro-modeling technique to approximate expensive functions and to replicate input–output relationships. Surrogate models are also called meta- or response surface models and posses beneficial qualities, such as cheaper evaluation costs and sometimes, better accuracy. To approximate complex simulation systems, deep learning based surrogate models are appropriate, as they are suitable to represent non-linear relationships and can handle high dimensional problems well. In the field of traffic forecasting, Vlahogianni [2015] introduced the concept of surrogate modelling and suggests some approaches to specific optimization problems that a transportation researcher may face when dealing with forecasting models. The approaches can be adapted to various transportation problems and the results convey that surrogate models can generate more accurate predictions in significantly reduced times.
Jariyasunant et al. [2014] developed a computation surrogate for travel feedback. They state that such surrogates can help to overcome scaling issues. Focusing on dynamic user equilibrium, Ge et al. [2020] formulated a dynamic user equilibrium traffic assignment model incorporating surrogate dynamic network loading models. They conclude that the usage of surrogate approaches reduces the computational times by up to 90%, while at the same time, leading to lower errors. Surrogate models are also used to assign Dedicated Bus Lanes (DBLs) across a large-scale road network. The results of numerical studies from Li et al. [2022] show that surrogate based methods for optimal DBL allocation scheme can save a substantial amount of time, especially with limited computational resource, while also leading to around 5% better network performance than the costly simulation approaches.
The concept of GNN based surrogate modelling in the context of travel demand modelling emerged very recently, with the pioneering work of Narayanan et al. [2024]. They laid a foundational groundwork by investigating the feasibility of developing a GCN based surrogate for strategic transport planning. Their study delve into both classification and regression setups, and based their analysis by generating data from Greater Munich metropolitan region. The experimental results show that GNNs have the potential to act as transport planning surrogates and the deeper models perform better than their shallow counterparts. However, the models suffer performance degradation with an increase in network size.
@@ 2.3 Summary and research gaps
The study of Narayanan et al. [2024] demonstrated the efficacy of GCNs and showcased how GCNs outperformed traditional machine learning models by effectively leveraging the graph structure inherent in transportation networks. Nevertheless, the limitations of GCNs (e.g., the neighbourhood aggregation function used in the GCN puts equal weight on all neighbours, while all neighbours may not be equally relevant) and that of the classification and regression setups necessitate the investigation of the use of advanced GNN architectures and the analysis of a new modelling setup, which will be the focus of this study.
@@ 3 Methodological framework
This study builds over the methodological framework established in Narayanan et al. [2024], as shown in Figure 1. Therefore, the procedural sequence is as follows: (i) data generation for the four-step model, (ii) simulation in the four-step model, (iii) data transformation (i.e., direct graph generation) and the creation of the training data (iv) GNN model training and model evaluation. The regularization methods (i.e., Dropout and GraphNorm), loss functions (cross entropy for the classification tasks and mean squared error for the regression tasks) and evaluation metrics (accuracy and F1-score for the classification tasks and mean absolute error and R2) from there are also utilised in this study. For detailed information of the setup, the reader is referred to the original study.
Figure 1: Methodological framework
As an addition to the existing setup, this study develops and investigates further GNN model arcitectures and additional data generation process, which are introduced in the following sections. Specifically, we describe the newly proposed GATv3 model, synthetic data generation and fine grained classification approach with the ability to obtain precise values as in the case of classical regression. Narayanan et al. [2024] implemented augmented data generation based on the network of Munich metropolitan region, but this study utilises the augmented data as well as synthetic data. With regards to the difference between the two, purely synthetic data is the one generated procedurally, whilst augmented implies that a true data source is employed, but modified to produce more data.
The implementation of the entire pipeline is performed in Python. The baseline models are estimated using Scikit-Learn [Pedregosa et al., 2011]. GNN models are implemented in PyTorch Geometric [Fey & Lenssen, 2019], which is an extension to the PyTorch library [Paszke et al., 2019]. The entire code (including the ones for data generation and processing) is provided in a private GitHub repository, which is found at https://github.com/nikita68/4_step_model_surrogate. Access to it is possible upon request. Nevertheless, the processed datasets are separately available at https://github.com/nikita6187/TransportPlanningDataset.
3.1 Graph attention networks
Narayanan et al. [2024] developed a GNN surrogate based on Graph Convolutions Networks (GCN). The neighbourhood aggregation function used in the GCN puts equal weight on all neighbours. However, not all neighbours may be equally relevant. An example would be an incoming road having a high traffic demand, whilst another having zero at a junction. Thus, the information relevant for the outgoing road comes more from the high demand incoming road. Therefore, in this section, the graph attention networks are presented, together with an extension to enable them to generalize to deeper networks.
GATv2 still suffer from the over-smoothing problem, much like the original GCN. This results in both models performing poorly when trying to stack more than 5 layers. To the best of the authors’ knowledge, there is no solution available to this problem at the time of this study. Inspired by the results of GCNII, an extension is proposed to use the same residual trick from Chen et al. [2020], which adds states from a lower layer directly to a higher layer and avoids the distinct local features getting vanished in the higher layers. Thus, the GATv2 formulation becomes as follows:
[…]
where,
$\alpha$ is a fixed hyperparameter
$\sigma$ is the activation function
$k$ represents all the direct neighbours of node $i$
$N_{ff}$ is a one layer MLP
[…]
[;][;][ ; ] operator concatenates the input vectors
Residual connections were originally proposed in the field of computer vision to mitigate vanishing gradient problems. In the current case, even though it is a minor change performed on the GATv2 model, it is an effective solution. It enables the possibility of developing deeper GAT models, as will be shown in later sections. From now, this model is denoted as $GATv3$. Note that, due to the extra computations, the number of GAT layers which can fit onto one GPU is significantly less than those of GCNs.
@@ 3.2 Synthetic data generation
For data generation, Narayanan et al. [2024] implemented augmented data generation, wherein the network of Munich metropolitan region is used to generate multiple subnetworks for model training. It is often seen that more data significantly helps deep learning models [Wang et al., 2017]. However, the original network data source utilised for the augmented dataset in Narayanan et al. [2024] has only a finite number of possible subnetworks, until the training dataset has a critical overlap with the validation and test datasets. To overcome this issue, while still providing arbitrarily more data, this study introduces an additional data generation procedure, namely the synthetic data generation.
A synthetic region generation algorithm is implemented, which does not use any original network as a source. This algorithm is motivated based on the works of Parish & Müller [2001], Smelik et al. [2014], Cogo et al. [2019]. Assuming that the target amount of nodes ($\mathds{G}_{nodes}$) is provided as input, along with the components for generating zones and zonal information (as identical to that of the augmented data generation procedure), the algorithm performs the following steps:
@@
The synthetic dataset is generated with 10000 samples, with a target number of nodes between 15 and 80. Note, the synthetic dataset is used in this paper only as supplementary training data for the augmented dataset with further details being provided in the experiments section (Section 3.3).
3.3 Experimental setup
This section will focus on two distinct setups, i.e., (i) an initial classification setup to compare the different GNNs and (iii) a fine-grained classification setup to exploit the stability of classification, which also includes the proposal of a small transformation procedure to convert the classification results into actual traffic values. Finally, the impact of additional data will be explored using the fine-grained classification setup.
3.3.1 classification task (comparison between GCNII, GATv2 and GATv3)